OpenStudy (anonymous):

If y varies inversely as x, and y = 18 when x = 2, find y when x = 4. Solve, using the quadratic formula: 2x2 - 3x + 1 = 0 Factor: 36x2 - 25

OpenStudy (anonymous):

#2 : 3 +/- sqrt(9 - (4(2)(-1)) / 2(2) 3 +/- sqrt(9 +8) / 4 3 +/- sqrt(17) / 4 [3 + sqrt(17)] / 4 and [3 - sqrt(17)] / 4 # 3 (6x - 5) (6x + 5) I Think

OpenStudy (anonymous):

#2 ... Doesn't have those answers?? @Reppinthe80

OpenStudy (anonymous):

o.o Really ?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

{1, 0.5} {2.4, -0.9} {-1, -0.5} {-2.4, 0.9} here are the answers @Reppinthe80

OpenStudy (anonymous):

These are 3 different questions?

OpenStudy (anonymous):

yes @DoYourHomework

OpenStudy (anonymous):

-_____- this Makes No Sense !!! O:

OpenStudy (anonymous):

That is why I need help lol... It's hard!! @Reppinthe80

OpenStudy (anonymous):

Ok...for the first one. When x is multiplied by 2, y is divided by 2 because it is inversely proportional so. x=2*2 = 4 So y=18/2 = 9 \[x = 4\] y = 9 see?

OpenStudy (anonymous):

Yessss.!! :)

OpenStudy (anonymous):

The second one you just have to factor This is the answer. Try to figure out how I got it. \[(x-1)(2x-1)\] x = 1 and x = 1/2 The third one is the same thing You just have to think a little bit. "What is the square root of 36 and 25?" \[(6x-5)(6x+5)\] Notice that when we FOIL this out. \[(6x*6x) + (6x*5 +6x*(-5)) -25\] (6x*5 +6x*(-5)) = 0 So you're back with your original \[36x^2 - 25\]

OpenStudy (anonymous):

Does this make sense? I hope I made it clear.